OpenStudy (anonymous):

Ally is wearing a witch’s costume for Halloween. Her hat is in the shape of a cone with a diameter of 7 inches and a height of 12 inches. What is the volume of Ally’s hat? Use 22/7 for pi.

OpenStudy (lgbasallote):

v = (pi hr^2)/3 v = [22(12)(7/2)^2]/7(3) v = 264(49/4)/21 v = (66)(49)/21 v = (22)(49)/7 v = 22(7) v = 154 cubic inches

OpenStudy (anonymous):

thanksss